# Perfect Divisibility and Coloring of Some Bull-Free Graphs

### v1

@@ Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edge, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [The Electron. J. of Combin. 29 (2022) P3.19.] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. Let $H\in${{odd balloon}, $\{P_{8},C_{5}\}\}$. In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible. We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph. As corollaries, these graphs is $\binom{\omega+1}{2}-$colorable. Notice that every odd torch contains an odd hole, a $P_{5}$, and a fork. Therefore, our results generalize the results of these scholars. Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
@@
In this paper, all graphs are finite and simple. As usual, let $P_{k}$ and $C_{k}$ be a path and a cycle on $k$ vertices respectively. We say that a graph $G$ contains a graph $H$ if $H$ is isomorphic to an induced subgraph of $G$. A graph $G$ is $H$-free if it does not contain $H$. Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow\{1,2,...,k\}$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. nonadjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$. For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$.
@@
[7] A graph $G$ is perfect if and only if $G$ is (odd hole, odd antihole)-free.
Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$. A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$. (See Figure 1)
@@
A graph is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoáng in [12]. Chudnovsky and Sivaraman [9] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [17] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. In this paper, we prove the following theorems.
@@
It is easy to verify that every odd torch contains an odd hole, a $P_{5}$, a fork, and the graph $E$. Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd balloon, bull)-free graphs. Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al. [17]; and Theorem 1.3 generalizes the result of Chen and Xu [5].
Recall that we use $F$ to denote the Mycielski-Grötzsch graph. It is certain that $F$ is ($P_{6}$, bull)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$.). In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible. Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible. In this paper, we prove the following theorem.
@@
Let $G$ be a $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
@@
Let $G$ be a (odd torch, $C_{3}$)-free graph. Then $\chi(G)\leq 3$.
@@
We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
@@ 2 Notations and Preliminary Results
In this section, we will prove Theorem 1.2. Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if for any vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect. The following useful lemmas is important to our proof.
@@
[14] Every minimal nonperfectly divisible graph has no homogeneous set.
@@
Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$. We next will show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$. If $n=3$, then $n^{\prime}=2$. To avoid an induced bull on $\{v_{1},v_{3},y,x,v_{4}\}$, we have that $y\sim v_{4}$; but then $\{v_{1},v_{4},y,x,v_{2}\}$ induces a bull, a contradiction. Hence, $n\geq 4$, and thus $v_{n}\sim v_{2}$ and $v_{n-1}\sim v_{1}$.
@@
Let $G$ be a minimal nonperfect divisible bull-free graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
Proof. It is certain that $G$ is connected. Suppose to its contrary that there exists a vertex $v\in V(G)$ such that $G[M(v)]$ is perfect. But then, $G[\{v\}\cup M(v)]$ is perfect and $G[N(v)]$ has clique number less than $\omega(G)$, which contradicts to that $G$ is minimal nonperfectly divisible. Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect. Moreover, since $G$ is minimal nonperfectly divisible, $G$ has no homogeneous set by Lemma 3.1. And so, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2. This proves Lemma 3.4. $\square$
@@
Let $G$ be a minimal nonperfectly divisible bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
@@
Proof of Theorems 1.2: Suppose the Theorem 1.2 is not true. Let $G$ be a minimal nonperfectly divisible (odd torch, bull)-free graph. By Lemma 3.4, we have that $G$ is connected, and
@@
By (3), we have that $y$ is not complete to $V(C)$ as $G[N(y)]$ is perfect. Let $3\leq n\leq k$ be the minimum integer such that $y\not\sim v_{n}$. But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction. This completes the proof of Theorem 1.2. $\square$
@@
In this section, we will prove Theorems 1.4 and 1.3. The following useful lemmas is important to our proof.
@@
If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull. Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$. Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$. Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume $u\sim v_{2}$.
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done. So, we may suppose $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is locally perfect, $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
@@
Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
@@
The proof of following lemma is similar with the proof of Lemma 4.1.
@@
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. From now on, the subscript is modulo 7 in the proof of the Lemma. For $1\leq i\leq 7$, let
@@
[…]
@@
[…]
@@
For every $u\in N(v)$, $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
@@
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 4.2. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.2. $\square$
@@
Proof of Theorems 1.3 and 1.4: It suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible. Suppose not, and let $G$ be a minimal nonperfectly divisible ($H$, bull)-free graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$. By Lemmas 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect. By Lemmas 4.3 and 4.4, we have that $\omega(G)\geq 3$ because $G$ is nonperfectly divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Lemma 3.5, $G[M(v)]$ is odd antihole-free except $C_{5}$. By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole. Therefore, by Lemmas 4.1 and 4.2, $G$ has a homogeneous set, which contradicts with Lemma 3.1. This completes the proof of Theorems 1.3 and 1.4. $\square$
@@
In this section, we will prove Theorem 1.5. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
@@
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraph of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free; this implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
@@
More generally, let ${\cal G}_{1}$ be a hereditary class of graphs in which every trianlge-free graph is $k$-colorable, let ${\cal G}_{2}$ be the class of bull-free graphs, and let ${\cal G}={\cal G}_{1}\cap{\cal G}_{2}$. In [6], Chudnovsky et al. proved that $\cal G$} is polynomially $\chi$-bounded. Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$.
@@
  • Research of the first author was supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926. And research of the second author was supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No. YKJ202448).

### v3

@@ Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edges, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [The Electron. J. of Combin. 29 (2022) P3.19.] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. Let $H\in${{odd torch}, $\{P_{8},C_{5}\}\}$. In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible. We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph. As corollaries, these graphs are $\binom{\omega+1}{2}-$colorable. Notice that every odd torch contains an odd hole, an induced $P_{5}$, and an induced fork. Therefore, our results generalize their findings. Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
@@
In this paper, all graphs are finite and simple. We follow [1] for undefined notations and terminology. As usual, let $P_{k}$ and $C_{k}$ be a path and a cycle on $k$ vertices respectively. We say that a graph $G$ contains a graph $H$ if $H$ is isomorphic to an induced subgraph of $G$. A graph $G$ is $H$-free if it does not contain $H$. Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$. A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow[k]$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. non-adjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$.
@@
[7] A graph $G$ is perfect if and only if $G$ is ((odd hole, odd antihole))-free.
Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$. A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$. (See Figure 1)
@@
A graph is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoáng in [12]. Chudnovsky and Sivaraman [9] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [15] proved that every (fork, bull)-free graph is perfectly divisible, and this class of graphs was generalized by Wu and Xu [19]. Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. In this paper, we prove the following theorems.
@@
It is easy to verify that every odd torch contains an odd hole, an induced $P_{5}$, an induced fork, and the graph $E$. Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd torch, bull)-free graphs. Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al. [15]; and Theorem 1.3 generalizes the result of Chen and Xu [5].
Recall that we use $F$ to denote the Mycielski-Grötzsch graph. It is certain that $F$ is ($P_{6}$, bull)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$.). In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible. Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible ? In this paper, we prove the following theorem.
@@
Let $G$ be an $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
@@
Let $G$ be an ((odd torch, $C_{3}$))-free graph. Then $\chi(G)\leq 3$.
@@
Let $G$ be a graph and let $w$ be a weight function on $V(G)$. We use $\omega_{w}(G)$ to denote the maximum weight of a clique in $G$. A graph $G$ is perfectly weight divisible [9, 12] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$, $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$. It is certain that if a graph $G$ is perfectly weight divisible, then $G$ is perfectly divisible. If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short).
As mentioned above, the structure and coloring problem of bull-free graphs has garnered significant attention, with new methods and research content [4, 6, 11, 13, 18] being continuously proposed. This paper utilizes the concept of perfectly weight divisibility to further analyze the structure of bull-free graphs and bull-free MNWD graphs, thereby deriving our main results. We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
@@ 2 Notations and Preliminary Results
Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if every vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect.
In this section, we will prove Theorem 1.2. The following useful lemmas is important to our proof.
@@
[9] Every MNWD graph has no homogeneous set.
@@
Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$. We next show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$. If $n=3$, then $n^{\prime}=2$. To avoid an induced bull on $\{v_{1},v_{3},y,x,v_{4}\}$, we have that $y\sim v_{4}$; but then $\{v_{1},v_{4},y,x,v_{2}\}$ induces a bull, a contradiction. Hence, $n\geq 4$, and thus $v_{n}\sim v_{2}$ and $v_{n-1}\sim v_{1}$.
@@
Let $G$ be a bull-free MNWD graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
Proof. It is certain that $G$ is connected. Since $G$ is MNWD, we have that for every vertex $v\in V(G)$, $G[M(v)]$ cannot be perfect as otherwise, $(\{v\}\cup M(v),N(v))$ is a partition of $V(G)$ such that for all positive integer weight function on $V(G)$, the $G[\{v\}\cup M(v)]$ is a perfect graph and $G[N(v)]$ has a smaller maximum weight of a clique than $G$, which contradicts with the minimality of $G$. Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect. Moreover, since $G$ is MNWD, $G$ has no homogeneous set by Lemma 3.1. Therefore, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2. This proves Lemma 3.4. $\square$
@@
Let $G$ be a bull-free MNWD graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
@@
Proof of Theorems 1.2: It suffices to show that each (odd torch, bull)-free graph is perfectly weight divisible. Suppose it is not true, and let $G$ be a (odd torch, bull)-free MNWD graph. By Lemma 3.4, we have that $G$ is connected, and
@@
By (3), we have that $y$ is not complete to $V(C)$ as $G[N(y)]$ is perfect. Let $3\leq n\leq k$ be the minimum integer such that $y\not\sim v_{n}$. But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction. Therefore, each (odd torch, bull)-free graph is perfectly weight divisible. This completes the proof of Theorem 1.2. $\square$
@@
In this section, we will prove Theorems 1.3 and 1.4. The following useful lemmas are important to our proof.
Lemma 4.1
Let $G$ be a graph with $\chi(G)\leq 3$. Then $G$ is perfectly weight divisible.
Proof. Let $w$ be a positive integer weight function on $V(G)$, and let $G^{\prime}$ be an induced subgraph of $G$. Since $\chi(G^{\prime})\leq 3$, there is a partition $(A,B)$ of $V(G^{\prime})$ such that $G^{\prime}[A]$ is bipartite and $B$ is a stable set of $G^{\prime}$. Let $X\subseteq B$ such that each vertex in $X$ is a isolated vertex of $G^{\prime}$. Then $(A\cup X,B\setminus X)$ is a partition of $G^{\prime}$ such that $G[A\cup X]$ is perfect, and $\omega_{w}(G[B\setminus X])<\omega_{w}(G^{\prime})$. This proves the lemma. $\square$
@@
If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull. Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$. Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$. Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume that $u\sim v_{2}$.
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume that $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done. So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is locally perfect, we have that $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
@@
Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
@@
The proof of the following lemma is similar to that of Lemma 4.2.
@@
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. For $1\leq i\leq 7$, let
@@
Suppose that $u$ is anticomplete to $\{v_{2},v_{7}\}$. Then $u$ has a neighbor in $\{v_{3},v_{4},v_{5},v_{6}\}$. Assume that $u$ is anticomplete to $\{v_{3},v_{6}\}$. If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, and we are done by Lemma 4.2. So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$, a contradiction. Therefore, $u$ is not anticomplete to $\{v_{3},v_{6}\}$ and we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{3},v_{4},u,v_{2},v_{6}\}$ induces a bull. So, $u\not\sim v_{6}$. To avoid a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$, we have that $u\not\sim v_{5}$. Moreover, to avoid an induced bull on $\{v_{3},v_{4},u,v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$. Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, then $u\in Y_{1}$ and we are done. So, $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
@@
If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{v_{5},v_{6},u,v_{4},v_{7}\}$ induces a bull. So $u$ has at most one neighbor in $\{v_{5},v_{6}\}$. If $u\sim v_{5}$, then there is a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$. Hence, $u$ is anticomplete to $\{v_{5},v_{6}\}$ by symmetry. And thus $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$. This proves Claim 4.2. $\square$
@@
For every $u\in N(v)$, we have that $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
@@
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 4.2. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.3. $\square$
@@
Proof of Theorems 1.3 and 1.4: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible. We will actually prove the following stronger conclusion: every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly weight divisible. Suppose not, and let $G$ be a ($H$, bull)-free MNWD graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$. By Lemma 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect. By Lemmas 4.14.4 and 4.5, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Lemma 3.5, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole. Therefore, by Lemmas 4.2 and 4.3, $G$ has a homogeneous set, which contradicts with Lemma 3.1. This completes the proof of Theorems 1.3 and 1.4. $\square$
@@
In this section, we will prove Theorem 1.5. A class of graphs is hereditary if it is closed under taking induced subgraphs. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
@@
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraphs of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free, which implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
@@
More generally, let ${\cal G}_{1}$ be a hereditary class of graphs in which every trianlge-free graph is $k$-colorable, let ${\cal G}_{2}$ be the class of bull-free graphs, and let ${\cal G}={\cal G}_{1}\cap{\cal G}_{2}$. In [6], Chudnovsky et al. proved that $\cal G$} is polynomially $\chi$-bounded. Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$ ?
@@
  • Ran Chen is supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926. Di Wu is supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No. YKJ202448).